# Hierarchical Mamba Meets Hyperbolic Geometry: A New Paradigm for Structured Language Embeddings

### v2

@@ 4 Experiments
We compare our proposed HiM models with their Euclidean counterparts, evaluated across four ontology datasets (i.e., DOID, FoodOn, WordNet, and SNOMED-CT) varying in scale and hierarchical complexity. [^2] Datasets are available from https://zenodo.org/records/14036213 and see Table C4 in Appendix C for details. (1) DOID offers a structured representation of human diseases through “is-a” relationships [32]. (2) FoodOn is a detailed ontology that standardizes food-related terminology, covering ingredients, dishes, and processes for nutritional classification and dietary research. It uses a hierarchical structure and borrows from existing ontologies like LanguaL [33]. (3) WordNet is a well-known benchmark that organizes English nouns, verbs, and adjectives into synonym sets connected by hypernym-hyponym relationships [34]. (4) SNOMED Clinical Terms (SNOMED-CT) is a comprehensive clinical terminology system used in electronic health records (EHRs). It organizes concepts (e.g., diagnoses, procedures, symptoms) into multiple hierarchies, linked by “is-a” and attribute relationships [35]. All datasets are derived from structured taxonomies and can be represented as directed acyclic graphs, where nodes denote entities and edges denote direct subsumption (i.e., parent-child) relations.
@@
We use 4 NVIDIA A100 GPUs with 80GB of memory each, distributed across a single compute node. Our model is implemented using the mamba-ssm library [10]. To define and operate over hyperbolic manifolds, we use GeoOpt [36], while DeepOnto [37] is employed to process and manage hierarchical structures in the ontology datasets. We leverage distributed data-parallel training with PyTorch’s DistributedDataParallel wrapper [38]. Our models were trained for ten epochs using the AdamW optimizer with a linear warm-up learning rate over the first 100 steps (target learning rate set to $41e-$ 4), and weight decay of $31e-$ 3. The linear warm-up is followed by a constant learning rate $41e-$ 4. The maximum gradient norm is clipped to 1.0. We employ a combination of hyperbolic clustering loss and hyperbolic centripetal loss during pretraining, with weights of 1.0 and 1.0, respectively. Our model incorporates several learnable parameters, such as scaling factor $\gamma$ (initialized to 0.01), curvature $c$ (initialized to 1). We implement dynamic margin parameters for losses $\alpha$ and $\beta$, which depend on the updated curvature. We use a batch size of 256 per GPU. To regularize the model during training, a dropout rate of 0.2 is applied following each Mamba2 block. The detailed train/validation/test splits for mixed-hop prediction and multi-hop inference tasks, can be found from Table C4 in Appendix C.
@@
We evaluated our HiM models on two key tasks designed to assess its hierarchical reasoning capabilities in ontology completion and knowledge graph inference: (1) multi-hop inference, which involves predicting the existence of indirect relationships (e.g., “dog is a vertebrate”) through transitive reasoning. (2) mixed-hop prediction, which focuses on estimating hierarchical distances between entities (e.g., 1-hop vs. 2-hop relations). Both tasks are formulated as classification problems based on hyperbolic distances. Detailed formulations are provided in Appendix D. We use three metrics for evaluation: F1 score, Precision, and Recall. Among them, F1 score serves as the primary metric, as it provides a balanced measure of precision and recall, which is critical for hierarchical reasoning tasks. Following prior work on these datasets [16], we exclude Accuracy due to its vulnerability to class imbalance, where negative samples significantly outnumber positive ones. During training, models are optimized using entity triplets (anchor, positive, negative) under a contrastive learning framework; however, evaluation is performed on entity pairs to directly assess subsumption prediction performance.
@@ 5 Results
We compare our proposed HiM models–HiM-Poincaré and HiM-Lorentz– against two Euclidean baselines (pretrained SentenceMamba-16M and finetuned SentenceMamba-16M) on four hierarchical datasets for two main downstream tasks: mixed-hop prediction and multi-hop inference. The pretrained SentenceMamba-16M is obtained by training on the SNLI dataset [29], while the fine-tuned SentenceMamba-16M model is initialized using Kaiming normal initialization for weights and zero initialization for biases, then fine-tuned on hierarchical datasets. Our HiM models share the SentenceMamba-16M backbone ($\approx 16$M parameters), but incorporate learnable curvature and are trained in hyperbolic space using both Poincaré and Lorentzian manifolds.
@@ 5.1 Comparison between HiM models and their Euclidean baselines
Table 1 presents a comprehensive performance comparison between HiM-Poincaré and HiM-Lorentz (SentenceMamba16M random initialized and fine-tuned for hierarchical datasets in hyperbolic manifolds), and their Euclidean baselines: the pretrained SentenceMamba-16M and fine-tuned SentenceMamba16M (random initialized but fine-tuned for our hierarchical datasets in Euclidean space). The two HiM models were trained with learnable curvature parameter $\mathcal{K}=-1/c$. A deeper curvature (smaller radius $r$ => smaller $c$ => larger/deeper curvature $\mathcal{K}$) allows us to exploit the hierarchical structure of the Hyperbolic manifold much better, as the hyperbolic embeddings are confined in the conical manifold compact within a smaller radius. The average $\delta$-hyperbolicity [39] for each dataset measures the tree-likeness of the graph by calculating the maximum deviation from the four-point condition. Values closer to 0 indicate a more hierarchical structure [40], making these datasets well-suited for hyperbolic embeddings. The corresponding $\delta$-hyperbolicity scores for the four datasets are reported in Table 1, reflecting a descending order of hierarchy complexity: DOID $\rightarrow$ SNOMED-CT $\rightarrow$ WordNet $\rightarrow$ FoodOn. The experimental results illustrate that HiM-Lorentz model achieves more robust and stable performance (with extremely small variance) in terms of the F1, precision, and recall values for both mixed-hop prediction and multi-hop inference tasks across four datasets. Moreover, HiM-Lorentz outperformed the HiM-Poincaré variant on the multi-hop inference task for both the WordNet and SNOMED-CT datasets, both are relatively large datasets and exhibit deeper hierarchies characterized by small $\delta$-hyperbolicity. However, in the case of FoodOn—which also has higher hyperbolicity—the Poincaré-based model achieved better performance.
5.2 Visualization of Hyperbolic Embeddings
Figure 2 presents the visualization of hyperbolic embeddings learned by HiM on the WordNet dataset, illustrating a representative hierarchical path: sport $\rightarrow$ skating $\rightarrow$ skateboarding. The HiM-trained embeddings exhibit tight clustering of semantically related nodes (e.g., skating and sport) in hyperbolic space, indicating enhanced semantic alignment. Moreover, the embeddings clearly capture the hierarchical structure: higher-level concepts such as sport are positioned closer to the origin, while more specific concepts like skateboarding are embedded farther from the origin in a compact and organized manner.
Table 1: Performance comparison of Pretrained, Finetuned across Euclidean manifold, and HiM models across Hyperbolic manifolds on various datasets (with varying average $\delta$-hyperbolicity). Pretrained SentenceMamba-16M is trained on SNLI; Finetuned and HiM load the pretrained model, but randomly initialize it before training. HiM uses learnable curvature for hyperbolic projections. The mean and standard deviation of F1, Precision and Recall scores were computed over five independent runs for each setting. (See details in Appendix E)
@@ 5.2 Visualization of Hyperbolic Embeddings
5.3 Hierarchical semantics encoded by hyperbolic geometry
To provide more interpretable results of our HiM models for the hierarchical learning, we conducted a deeper geometric analysis of the hyperbolic entity embeddings for semantically related WordNet entities under HiM-Poincaré and HiM-Lorentz manifolds (see Fig. 2 for the visualization of hyperbolic embeddings). Specifically, we computed three key metrics with the learned hyperbolic embeddings: 1) “hyperbolic geodesic distances” between each pair of entities, 2) “h-norm” represents the norm distance from the origin, a higher h-norm often indicates a deeper or more specific concept in the hierarchy, 3) “depth” is the WordNet tree depth. In both sets of entities, the h-norm correlates strongly with the hierarchical depth, see Fig. 3. For instance, in Table 2, the entity sport (depth 9) has an h-norm of 0.55, while skateboarding (depth 12) has an h-norm of 2.25, reflecting the hierarchical expansion from general to specific concepts. However, a key difference emerges when comparing the two manifolds: HiM-Lorentz consistently produces smaller hyperbolic distances and h-norms compared to HiM-Poincaré. For example, in Tables 3, the parent-child relationships (e.g., sport $\rightarrow$ skating $\rightarrow$ skateboarding) embedded with HiM-Lorentz model exhibit tighter distances and h-norm gradients compared to those obtained by HiM-Poincaré, as shown in Table 2.
This reduction in distances under the Lorentz manifold is advantageous for hierarchical modeling. The Lorentz model’s unbounded nature avoids the boundary constraints of the Poincaré ball, which can lead to numerical instability near the boundary. By mapping embeddings into an unbounded hyperboloid, HiM-Lorentz achieves tighter clustering of related entities (e.g., sport and skating: 0.73 in Lorentz vs. 1.67 in Poincaré) while maintaining the hierarchical structure. This tighter clustering enhances the model’s ability to distinguish fine-grained relationships, especially in deeper hierarchies, as evidenced by the smaller standard deviations of HiM-Lorentz in performance metrics (Table 1).
@@ 5.3 Hierarchical semantics encoded by hyperbolic geometry
Figure 2: Visualization of HiM’s embeddings trained on the WordNet dataset in the Poincaré ball manifold. Left: The full hyperbolic space, illustrating the distribution of entities with parent nodes positioned closer to the origin and child nodes extending toward the boundary, reflecting the exponential expansion of hyperbolic geometry. Right: A zoomed-in view emphasizing the hierarchical structure, such as the path sport $\to$ skating $\to$ skateboarding. Dots represent the entities, with colors indicating hierarchical relationships. For a selected node “skating” denoted by green dot, the blue node denotes its parent nodes (e.g., sport), and red indicates its hard negatives, such as siblings/cousins (e.g., rowing). Yellow nodes (e.g, skateboarding, speed skating) indicate children nodes of the selected node (skating), meaning the grandchildren nodes of the blue node (sport).
Table 2: Hyperbolic distances, h-norms, and depths for sports-related entities (from Fig. 2) using HiM-Poincaré, sorted by increasing depth.
sport outdoor sport skating rowing skateboarding speed skating
sport 0.00 1.17 1.67 1.62 2.43 2.40
outdoor sport 1.17 0.00 1.90 1.95 2.66 2.62
skating 1.67 1.90 0.00 2.36 3.12 3.07
rowing 1.62 1.95 2.36 0.00 3.10 3.11
skateboarding 2.43 2.66 3.12 3.10 0.00 3.76
speed skating 2.40 2.62 3.07 3.11 3.76 0.00
h-norm 0.55 0.97 1.48 1.48 2.25 2.25
depth 9 10 11 11 12 12
Table 3: Hyperbolic distances, h-norms, and depths for sports-related entities (from Fig. 2) using HiM-Lorentz, sorted by increasing depth.
sport outdoor sport skating rowing skateboarding speed skating
sport 0.00 0.55 0.73 0.71 0.87 0.87
outdoor sport 0.55 0.00 0.79 0.87 0.98 0.95
skating 0.73 0.79 0.00 0.95 1.08 1.08
rowing 0.71 0.87 0.95 0.00 1.07 1.07
skateboarding 0.87 0.98 1.08 1.07 0.00 1.20
speed skating 0.87 0.95 1.08 1.07 1.20 0.00
h-norm 0.30 0.48 0.65 0.65 0.81 0.81
depth 9 10 11 11 12 12
@@ 6 Conclusion
Lorentz embeddings can provide a more natural fit for large-scale datasets with intricate hierarchical patterns compared to other geometries, potentially enhancing performance and interpretability. By demonstrating how a Lorentzian manifold can be effectively deployed for hyperbolic sentence representations, this paper aims to motivate further exploration of hyperbolic geometry in diverse real-world applications, ultimately broadening the scope and impact of geometry-aware neural architectures.
Limitations and Future Scope: HiM’s application to broader NLP scenarios, such as real-time inference, remains underexplored. Investigating HiM’s potential for efficient temporal dependency modeling in intricate long-range hierarchical classification tasks holds significant promise and study its practical applications. Future work could explore integrating CLIP-style pretraining to incorporate multimodal data (e.g., text and images) for tasks like visual question answering, potentially building on work such as Cobra [41], which demonstrates the potential of extending Mamba models for efficient multimodal language modeling.
Acknowledgments and Disclosure of Funding
We would like to acknowledge the funding support from the DOE SEA-CROGS project (DE-SC0023191), AFOSR project (FA9550-24-1-0231), and the Grace Hopper AI Research Institute (GHAIRI) Seed Grant at NJIT. We also thank the computing resources provided by the High Performance Computing (HPC) facility at NJIT.
Appendix
Appendix A Preliminaries
A.1 Hyperbolic Geometry
@@
[…]
This representation has gained attention in machine learning due to relatively simple re-parameterizations for gradient-based updates, thus facilitating the embedding of hierarchically structured data [4].
@@
where $\langle\cdot,\cdot\rangle_{\!M}$ denotes the Minkowski inner product, typically $-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$. The hyperbolic distance between two points $\mathbf{x}$ and $\mathbf{y}$ then appears in the form:
@@
Compared to the Poincaré ball, this approach can sidestep certain numerical instabilities near the boundary because vectors are not constrained to lie within a finite radius. Moreover, Lorentz-based formulations often allow more direct computation of geodesics and exponential maps, making them advantageous for large-scale hyperbolic embeddings [11]. Krioukov et al. [42] provide a theoretical foundation for the hyperbolic geometry of complex networks, showing that many real-world networks naturally embed into hyperbolic spaces, supporting our choice of the Poincaré and Lorentzian models for hierarchical language embeddings.
Figure A1: Illustration of word embeddings in Euclidean (Left) vs. Hyperbolic Spaces for hierarchical representation in Poincaré (Top right) and Lorentzian Manifolds (Bottom right).
The Fig. A1 illustrates how hierarchies are represented differently in Euclidean space, the Poincaré manifold, and the Lorentzian manifold. On the left, the hierarchical structure is arranged as a standard tree. While the relationships are maintained, Euclidean space does not naturally encode hierarchical distances in 2D. In Fig. A1 the upper-right diagram shows the hierarchy embedded into the Poincaré ball (the root/origin being at the center). The more generalized parent nodes are positioned near origin, and descendant nodes extend outward near the margin. This representation captures the exponential growth of hierarchical structures, where sibling nodes are placed far apart in terms of geodesic distance. The lower-right diagram visualizes the same hierarchy embedded in the Lorentz hyperboloid. The Lorentzian manifold in $\mathbb{R}^{n+1}$ consists of $n$ spatial dimensions and one time-like dimension ($x_{0}$). The origin is at the center-bottom of the Hyperboloid, and nodes are arranged along the hyperboloid surface. More generalized parent nodes are positioned near the bottom, and the descendants keep extending upward on the cone of Lorentz. Unlike the Poincaré model, which confines embeddings within a finite ball, the Lorentz model represents hierarchies in an unbounded space, making it particularly suitable for representing deeply nested hierarchies.
A.2 Mamba2
Mamba2 is a state-space model (SSM) introduced by Tri Dao and Albert Gu that refines the original Mamba architecture with improved performance and simplified design [10]. Mamba-2 builds upon the original Mamba architecture by introducing the State Space Duality (SSD) framework, which establishes theoretical connections between State Space Models (SSMs) and attention mechanisms. Mamba2 achieves 2-8$\times$ faster processing while maintaining competitive performance compared to Transformers for language modeling tasks. In order to formulate the overall computation for a single Mamba2 block, let $\mathbf{x}_{1:L}=[\mathbf{x}_{1},\dots,\mathbf{x}_{L}]$ be the token (or embedding) sequence for a given input. A single Mamba2 block transforms $\mathbf{x}_{1:L}$ into an output sequence $\mathbf{y}_{1:L}$. Each input token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is first normalized via RMSNorm. RMS normalization ensures that the norm of the embeddings remains stable across different inputs, preventing extreme values from causing instability during training.

### v3

@@ 4 Experiments
We compare our proposed HiM models with their Euclidean counterparts, evaluated across four ontology datasets (i.e., DOID, FoodOn, WordNet, and SNOMED-CT) varying in scale and hierarchical complexity. [^2] Datasets are available from https://zenodo.org/records/14036213 and see Table 3 in Appendix A.3 for details. (1) DOID offers a structured representation of human diseases through “is-a” relationships (Schriml et al., 2012). (2) FoodOn is a detailed ontology that standardizes food-related terminology, covering ingredients, dishes, and processes for nutritional classification and dietary research. It uses a hierarchical structure and borrows from existing ontologies like LanguaL (Dooley et al., 2018). (3) WordNet is a well-known benchmark that organizes English nouns, verbs, and adjectives into synonym sets connected by hypernym-hyponym relationships (Miller, 1995). (4) SNOMED Clinical Terms (SNOMED-CT) is a comprehensive clinical terminology system used in electronic health records (EHRs). It organizes concepts (e.g., diagnoses, procedures, symptoms) into multiple hierarchies, linked by “is-a” and attribute relationships (Stearns et al., 2001). All datasets are derived from structured taxonomies and can be represented as directed acyclic graphs, where nodes denote entities and edges denote direct subsumption (i.e., parent-child) relations.
@@
We use 4 NVIDIA A100 GPUs with 80GB of memory each, distributed across a single compute node. Our model is implemented using the mamba-ssm library (Dao & Gu, 2024). To define and operate over hyperbolic manifolds, we use GeoOpt (Kochurov et al., 2020), while DeepOnto (He et al., 2024a) is employed to process and manage hierarchical structures in the ontology datasets. We leverage distributed data-parallel training with PyTorch’s DistributedDataParallel wrapper (Paszke, 2019). Our models were trained for ten epochs using the AdamW optimizer with a linear warm-up learning rate over the first 100 steps (target learning rate set to 1e41e-4), and weight decay of 1e31e-3. The linear warm-up is followed by a constant learning rate 1e41e-4. The maximum gradient norm is clipped to 1.0. We employ a combination of hyperbolic clustering loss and hyperbolic centripetal loss during pretraining, with weights of 1.0 and 1.0, respectively. Our model incorporates several learnable parameters, such as scaling factor $\gamma$ (initialized to 0.01), curvature $c$ (initialized to 1). We implement dynamic margin parameters for losses $\alpha$ and $\beta$, which depend on the updated curvature. We use a batch size of 256 per GPU. To regularize the model during training, a dropout rate of 0.2 is applied following each Mamba2 block. The detailed train/validation/test splits for mixed-hop prediction and multi-hop inference tasks, can be found from Table 3 in Appendix A.3.
@@
We evaluated our HiM models on two key tasks designed to assess its hierarchical reasoning capabilities in ontology completion and knowledge graph inference: (1) multi-hop inference, which involves predicting the existence of indirect relationships (e.g., “dog is a vertebrate”) through transitive reasoning. (2) mixed-hop prediction, which focuses on estimating hierarchical distances between entities (e.g., 1-hop vs. 2-hop relations). Both tasks are formulated as classification problems based on hyperbolic distances. Detailed formulations are provided in Appendix A.4. We use three metrics for evaluation: F1 score, Precision, and Recall. Among them, the F1 score serves as the primary metric, as it provides a balanced measure of precision and recall, which is critical for hierarchical reasoning tasks. Following prior work on these datasets (He et al., 2024b), we exclude Accuracy due to its vulnerability to class imbalance, where negative samples significantly outnumber positive ones. During training, models are optimized using entity triplets (anchor, positive, negative) under a contrastive learning framework; however, evaluation is performed on entity pairs to directly assess subsumption prediction performance.
@@ 5 Results
We compare our proposed HiM models–HiM-Poincaré and HiM-Lorentz– against two Euclidean baselines (pretrained SentenceMamba-16M and finetuned SentenceMamba-16M) on four hierarchical datasets for two main downstream tasks: mixed-hop prediction and multi-hop inference. The pretrained SentenceMamba-16M is obtained by training on the SNLI dataset (Bowman et al., 2015), while the fine-tuned SentenceMamba-16M model is initialized using Kaiming normal initialization for weights and zero initialization for biases, then fine-tuned on hierarchical datasets. Our HiM models share the SentenceMamba-16M backbone ($\approx 16$M parameters), but incorporate learnable curvature and are trained in hyperbolic space using both Poincaré and Lorentzian manifolds.
@@ 5.1 Comparison between HiM models and their Euclidean baselines
We present a comprehensive comparison of HiM-Poincaré, HiM-Lorentz, and their Euclidean baselines, including the pretrained SentenceMamba-16M and the fine-tuned SentenceMamba-16M (randomly initialized and trained on hierarchical datasets in Euclidean space), in Table 1. Both HiM models were trained with learnable curvature parameter $\mathcal{K}=-1/c$. A deeper curvature (smaller radius $r$ => smaller $c$ => larger/deeper curvature $\mathcal{K}$) allows us to exploit the hierarchical structure of the Hyperbolic manifold much better, as the hyperbolic embeddings are confined in the conical manifold compact within a smaller radius. The average $\delta$-hyperbolicity (Gromov, 1987) for each dataset measures the tree-likeness of the graph by calculating the maximum deviation from the four-point condition. Values closer to 0 indicate a more hierarchical structure (Adcock et al., 2013), making these datasets well-suited for hyperbolic embeddings. The corresponding $\delta$-hyperbolicity scores for the four datasets are reported in Table 1, reflecting a descending order of hierarchy complexity: DOID $\rightarrow$ SNOMED-CT $\rightarrow$ WordNet $\rightarrow$ FoodOn. The experimental results illustrate that HiM-Lorentz model achieves more robust and stable performance (with extremely small variance) in terms of the F1, precision, and recall values for both mixed-hop prediction and multi-hop inference tasks across four datasets. Moreover, HiM-Lorentz outperformed the HiM-Poincaré variant on the multi-hop inference task for both the WordNet and SNOMED-CT datasets, both are relatively large datasets and exhibit deeper hierarchies characterized by small $\delta$-hyperbolicity. However, in the case of FoodOn—which also has higher hyperbolicity—the Poincaré-based model achieved better performance.
Table 1: Performance comparison of Pretrained, Finetuned across Euclidean manifold, and HiM models across Hyperbolic manifolds on various datasets (with varying average $\delta$-hyperbolicity). Pretrained SentenceMamba-16M is trained on SNLI; Finetuned and HiM load the pretrained model, but randomly initialize it before training. HiM uses learnable curvature for hyperbolic projections. The mean and standard deviation of F1, Precision and Recall scores were computed over five independent runs for each setting. (See details in Appendix A.6)
@@
5.2 Comparison with Hyperbolic Transformer Baseline
To further evaluate the effectiveness of HiM, we compare it with a hyperbolic transformer model (HiT) (He et al., 2024b), which was designed by fine tuning a pretrained model (i.e., all-MiniLM-L6-v2) with low curvature ($\mathcal{K}=-1/384$, nearly Euclidean). To ensure a fair comparison while preserving the hyperbolic nature of the space, we randomly initialize our HiM and evaluate it against a randomly initialized version of HiT (HiT*). We conducted ablations across two different curvatures ($\mathcal{K}=-1.0$ and 1/d-1/d where embedding dimension $d=384$) and model sizes (16M and 32M parameters) on both Poincaré and Lorentz manifolds. The results in Table 2 show that HiM consistently outperforms the transformer-based HiT* model across both manifolds and most experimental configurations.
Table 2: F1 scores comparing HiM models with HiT* across different parameter scales and curvature settings on mixed-hop prediction tasks. HiT* (Hyperbolic Transformer with random initialization) uses identical hyperbolic projections, loss functions, and manifolds as HiM, differing only in the use of Transformer architecture instead of Mamba. Results demonstrate consistent advantages of the Mamba architecture over Transformers in hyperbolic settings.
#Param. Curvature $(\mathcal{K})$ HiM-Poincaré HiT*-Poincaré HiM-Lorentz HiT*-Lorentz
WordNet DOID WordNet DOID WordNet DOID WordNet DOID
16M 1.0-1.0 0.859 0.902 0.846 0.837 0.850 0.890 0.838 0.844
16M 1/d-1/d 0.788 0.863 0.840 0.825 0.781 0.704 0.525 0.520
32M 1.0-1.0 0.849 0.883 0.836 0.831 0.837 0.869 0.829 0.832
32M 1/d-1/d 0.769 0.839 0.809 0.835 0.815 0.711 0.465 0.530
Notably, performance degrades significantly under low curvature settings ($\mathcal{K}=-1/d$), particularly in the Lorentz manifold, where HiT* shows substantial performance drops. This suggests that stronger hyperbolic curvature ($\mathcal{K}=-1.0$) is essential for effective hierarchical modeling. Under our fixed curvature, Poincaré’s bounded nature enables more stable norm dispersion and discriminative gradient flow, particularly when combined with variance regularization in our centripetal loss. In contrast, Lorentz embeddings tend to collapse toward the hyperboloid’s shell where time-like distances flatten and norm-based separation weakens. To provide broader context for our results, we include additional analysis with comparing our model to GPT-4o on zero-shot mixed-hop prediction and also study HiM’s computational efficiency in terms of sequence length in Appendix A.7.
5.3 Visualization of Hyperbolic Embeddings
To demonstrate how HiM captures hierarchical structure in the learned embeddings, we visualize the hyperbolic representations on a representative semantic hierarchy with WordNet. The hyperbolic embeddings learned by HiM is presented in Figure 2, which illustrates a representative hierarchical path, sport $\rightarrow$ skating $\rightarrow$ skateboarding. The HiM-trained embeddings exhibit tight clustering of semantically related nodes (e.g., skating and sport) in hyperbolic space, indicating enhanced semantic alignment. Moreover, the embeddings clearly capture the hierarchical structure, as higher-level concepts such as sport are positioned closer to the origin, while more specific concepts like skateboarding are embedded farther from the origin in a compact and organized manner. More details and geometric analysis of these hyperbolic embeddings, including quantitative metrics comparing h-norms, geodesic distances, and hierarchical depth correlations across both Poincaré and Lorentz manifolds, is provided in Appendix A.5.
@@ 5.3 Visualization of Hyperbolic Embeddings
Figure 2: Visualization of HiM’s embeddings trained on the WordNet dataset in the Poincaré ball manifold. Left: The full hyperbolic space, illustrating the distribution of entities with parent nodes positioned closer to the origin and child nodes extending toward the boundary, reflecting the exponential expansion of hyperbolic geometry. Right: A zoomed-in view emphasizing the hierarchical structure, such as the path sport $\to$ skating $\to$ skateboarding. Dots represent the entities, with colors indicating hierarchical relationships. For a selected node “skating” denoted by the green dot, the blue node denotes its parent nodes (e.g., sport), and red indicates its hard negatives, such as siblings/cousins (e.g., rowing). Yellow nodes (e.g, skateboarding, speed skating) indicate children nodes of the selected node (skating), meaning the grandchildren nodes of the blue node (sport).
@@ 6 Conclusion
Lorentz embeddings can provide a more natural fit for large‐scale datasets with intricate hierarchical patterns compared to other geometries, potentially enhancing performance and interpretability. By demonstrating how a Lorentzian manifold can be effectively deployed for hyperbolic sentence representations, this paper aims to motivate further exploration of hyperbolic geometry in diverse real‐world applications, ultimately broadening the scope and impact of geometry‐aware neural architectures. Investigating HiM’s potential for efficient temporal dependency modeling in intricate long-range hierarchical classification tasks holds significant promise and study its practical applications. Future work could explore integrating CLIP-style pretraining to incorporate multimodal data (e.g., text and images) for tasks like visual question answering, or potentially building on work such as Cobra (Zhao et al., 2025), which demonstrates the potential of extending Mamba models for efficient multimodal language modeling.
Acknowledgments
This work was supported by the DOE SEA-CROGS project (DE-SC0023191), AFOSR project (FA9550-24-1-0231). We also acknowledge the computing resources provided by the High Performance Computing (HPC) facility at NJIT.
Appendix A Appendix
A.1 Preliminaries
A.1.1 Hyperbolic Geometry
@@
$d_{\mathcal{P}}(\mathbf{x},\mathbf{y})=\sqrt{c}\cdot\mathrm{arcosh}\mathopen{\left(\,1+2\,\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{(1-\|\mathbf{x}\|^{2}/c)(1-\|\mathbf{y}\|^{2}/c)}\right)}.$ (11)
This representation has gained attention in machine learning due to relatively simple re-parameterizations for gradient-based updates, thus facilitating the embedding of hierarchically structured data (Nickel & Kiela, 2017).
@@
where $\langle\cdot,\cdot\rangle_{\!M}$ denotes the Minkowski inner product, typically $x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$. The hyperbolic distance between two points $\mathbf{x}$ and $\mathbf{y}$ then appears in the form:
@@
Compared to the Poincaré ball, this approach can sidestep certain numerical instabilities near the boundary because vectors are not constrained to lie within a finite radius. Moreover, Lorentz-based formulations often allow more direct computation of geodesics and exponential maps, making them advantageous for large-scale hyperbolic embeddings (Nickel & Kiela, 2018). Krioukov et al. (2010) provides a theoretical foundation for the hyperbolic geometry of complex networks, showing that many real-world networks naturally embed into hyperbolic spaces, supporting our choice of the Poincaré and Lorentzian models for hierarchical language embeddings.
Figure 3: Illustration of word embeddings in Euclidean (Left) vs. Hyperbolic Spaces for hierarchical representation in Poincaré (Top right) and Lorentzian Manifolds (Bottom right).
We can see how hierarchies are represented differently in Euclidean space, the Poincaré manifold, and the Lorentzian manifold as illustrated in Figure 3. On the left, the hierarchical structure is arranged as a standard tree. While the relationships are maintained, Euclidean space does not naturally encode hierarchical distances in 2D. In Figure 3 the upper-right diagram shows the hierarchy embedded into the Poincaré ball (the root/origin being at the center). The more generalized parent nodes are positioned near origin, and descendant nodes extend outward near the margin. This representation captures the exponential growth of hierarchical structures, where sibling nodes are placed far apart in terms of geodesic distance. The lower-right diagram visualizes the same hierarchy embedded in the Lorentz hyperboloid. The Lorentzian manifold in $\mathbb{R}^{n+1}$ consists of $n$ spatial dimensions and one time-like dimension ($x_{0}$). The origin is at the center-bottom of the Hyperboloid, and nodes are arranged along the hyperboloid surface. More generalized parent nodes are positioned near the bottom, and the descendants keep extending upward on the cone of Lorentz. Unlike the Poincaré model, which confines embeddings within a finite ball, the Lorentz model represents hierarchies in an unbounded space, making it particularly suitable for representing deeply nested hierarchies.
A.1.2 Mamba2
[…]